OpenStudy (anonymous):

Identify the GCF of 10x^4y^3 − 5x^3y^2 + 20x^2y.

jimthompson5910 (jim_thompson5910):

GCF of 10, -5, and 20 is 5 GCF of x^4y^3, x^3y^2, and x^2y is x^2y So the GCF of the three terms is 5x^2y

OpenStudy (anonymous):

eplain how to completely factor 2x2 − 8x − 42.

myininaya (myininaya):

find two factors of 2(-42) that have product 2(-42) and have sum -8; 2(-42)=4(-21)=12(-7)=6(-14) -8=-14+6 -8x=-14x+6x so replace -8x with -14x+6x and then factor by grouping